4x+3y=1, 2x+5y=11 solve by elimination by subtraction
Answer by mananth(16946)

4 x + 3 y = 1 .............1
2 x + 5 y = 11 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 3
-20 x -15 y = -5
6 x 15 y = 33
Add the two equations
-14 x = 28
/ -14
x = -2
plug value of x in (1)
4 x + 3 y = 1
-8 + 3 y = 1
3 y = 1 + 8
3 y = 9
y = 3